OpenStudy (toxicsugar22):

(x-4)^2/9-(y+2)^2/16=1 give the exact coordinates of any four points on the hyperbola. and show me how you get this

OpenStudy (anonymous):

rework the equation to be an x= or y= equation and go about it that way. you can put in any point for the variable remaining on the right side and solve it for the missing one.

OpenStudy (toxicsugar22):

i tried can u help me durther please

OpenStudy (anonymous):

set up the equation so that the x term is on one side and the y term is on the left, then multiply through by the lowest common demoninator in order to eliminate the fractions. what do you get?

OpenStudy (toxicsugar22):

can u helo me i dont know how to do this and if u help me with this one i will know for future

OpenStudy (anonymous):

move the y term to the right side, which gives you (x-4)^2/9 = (y+2)^2/16 + 1. multiply both sides of the equation by 9 to eliminate the fraction. take the square root of both sides in order to eliminate the exponential on the left. then add 4 to both sides in order to get x alone on the left. this all gives you\[x = 4 + \sqrt{9(y+2)^2)/16 + 9}\]

OpenStudy (anonymous):

then all you have to do is plug in any 4 numbers for x and get the corresponding y points by solving.